OpenStudy (anonymous):

The function f(x) =4x^4-2x^3+1 has a point of inflection at x=?

OpenStudy (mathstudent55):

Find the second derivative of the function. Set it equal to zero. Solve for x.

OpenStudy (anonymous):

Any work on this ??

OpenStudy (anonymous):

I found the second derivative, 48x^2-12x. But im not sure how to solve it when I set x=0

OpenStudy (campbell_st):

well you set the 2nd derivative equal to zero... and you'll have 2 points of infection so if \[\frac{d^2y}{dx^2} = 48x^2 - 12x\] then let\[\frac{d^2y}{dx^2} = 0\] so solve \[0 = 48x^2 - 12x\]

OpenStudy (campbell_st):

one of the infection points... is a horizontal point of inflection

OpenStudy (mathstudent55):

Factor the right side.